# Structural-Spectral Graph Convolution with Evidential Edge Learning for Hyperspectral Image Clustering

### v2

@@ I Introduction
The second limitation is the inaccurate superpixel topological graph. Typically, the superpixel topological graph is constructed based on either spatial adjacency or spectral similarity. The former connects spatially adjacent superpixels as neighbors [25, 38], while the latter determines neighbors based on spectral similarity [30, 32, 33]. Since the graph is constructed without supervision, both types of graphs inevitably contain incorrect edges that connect two superpixel nodes of different classes. Accordingly, these incorrect edges bring semantically different neighbors, leading to distorted node representations and confused class semantics during GNN aggregation. Fig. 1b illustrates the impact of deleting incorrect edges on clustering results. As the deletion ratio increases, clustering performance gradually improves, highlighting the importance of dealing with incorrect edges in the graph.
@@ III-D Training Procedure
We illustrate the training procedure of the proposed method in algorithm 1. In each epoch, we first compute the empirical edge weights and the predicted edge weights to update the adjacency matrix. Second, we perform forward-propagation, computing node representations, prototypes and loss functions. Third, network parameters are optimized. Fourth, under evaluation mode, we apply spherical K-means on the output of $f^{\prime}(\cdot)$ to get the current superpixel clustering assignments and then update the prototypes. Finally, we transform the clustering result from the superpixel level to the pixel level by assigning each superpixel’s label to its internal pixels.
@@
5:  Randomly sample pixels to get $\{\bm{x}^{+}\}$.
6:  Forward $f(\bm{x}),f^{\prime}(\bm{x}^{+})$ by Eq. (12).
7:  Compute $\mathcal{L}_{NA}$ by Eq. (7).
8:  Compute $\{\bm{\mu}\},\{\bm{\mu}^{+}\}$ by Eq. (8).
9:  Compute $\mathcal{L}_{PC}$ by Eq. (9).
10:  Compute $\mathcal{L}$ by Eq. (22).
11:  Update $f(\cdot)$ with SGD optimizer.
12:  Update $f^{\prime}(\cdot)$ by Eq. (11).
13:  Evaluation mode, forward $f^{\prime}(\bm{x})$, perform K-means to get $\mathbf{R}$, and compute $\{\bm{\mu}\}$ by Eq. (8).
14:end for
15: Transform $\mathbf{R}$ from the superpixel level to the pixel level.
16:return $\mathbf{R}$
@@ IV-A Datasets
TABLE II: Clustering Performance on HSI Datasets, Indian Pines, Pavia University, Botswana and Trento are Abbreviated to IP, PU, BO and TR, $\bullet$$\circ$ Indicates whether Our Method is Statistically Superior/Inferior to the Compared Method According to the Pairwise $t$-Test at 0.05 Significance Level.
@@
SSGCO-
EGAEL
IP ACC 34.72±1.00 47.99±0.46 53.48±0.84 58.29±1.19 57.10±3.39 65.41±3.37 64.70±0.00 50.58±4.46 67.40±1.17 55.88±0.00 70.01±1.95
$\mathcal{K}$ 28.39±1.10 42.13±0.57 48.54±1.12 54.54±1.22 52.18±3.56 61.10±3.50 60.24±0.00 44.01±4.27 63.62±1.34 49.59±0.00 66.55±2.12
NMI 43.24±0.57 57.81±0.44 59.47±1.40 63.65±0.35 63.68±1.03 67.65±1.65 71.97±0.00 50.21±1.80 67.54±1.26 54.03±0.00 71.17±1.31
ARI 21.02±0.99 31.46±0.55 36.13±1.42 40.21±0.49 45.35±4.75 48.99±3.76 52.93±0.00 31.51±3.83 56.02±1.94 36.51±0.00 55.99±2.07
Precision 37.74±1.15 43.93±1.72 42.74±2.88 51.88±1.62 50.59±1.81 51.34±1.87 59.63±0.00 45.34±2.71 52.66±2.19 49.61±0.00 58.72±2.23
Recall 40.27±0.38 47.42±5.06 51.12±0.62 57.11±1.97 49.79±3.89 51.11±3.16 59.08±0.00 46.37±1.38 61.34±3.48 52.39±0.00 65.77±4.42
F1 34.24±0.87 40.65±1.34 40.76±0.92 49.50±1.76 46.00±2.58 48.07±3.10 55.19±0.00 44.30±2.74 52.30±1.89 44.67±0.00 57.65±2.38
Purity 52.26±0.87 57.78±0.51 62.87±0.65 69.58±1.25 68.94±1.51 72.52±2.06 75.32±0.00 56.51±1.36 73.94±1.78 59.33±0.00 77.18±1.03
PU ACC 51.15±0.02 55.17±0.00 56.08±3.69 57.24±0.34 58.56±5.21 59.42±0.28 60.13±0.00 61.03±1.80 62.62±0.38 63.94±0.00 70.00±3.46
$\mathcal{K}$ 41.49±0.06 45.57±0.00 44.89±5.06 48.28±0.36 49.15±5.07 50.98±0.47 49.06±0.00 50.18±2.12 54.82±0.46 53.02±0.00 62.29±3.65
NMI 53.97±0.01 51.48±0.00 47.34±4.27 53.69±0.39 53.72±2.58 62.93±0.08 48.15±0.00 56.41±1.44 57.84±1.20 61.89±0.00 61.90±3.18
ARI 31.02±0.03 35.26±0.00 42.84±5.04 42.64±0.20 43.65±6.69 49.04±0.18 43.12±0.00 42.94±2.63 46.19±0.89 44.91±0.00 57.26±8.57
Precision 51.83±0.01 43.00±0.00 40.91±4.93 50.29±0.39 48.16±4.44 56.77±0.70 48.68±0.00 53.97±3.11 56.32±1.16 56.21±0.00 57.82±1.22
Recall 53.59±0.01 48.81±0.00 46.19±7.95 53.75±0.35 50.75±3.47 55.63±0.51 50.37±0.00 55.34±1.58 58.96±0.42 59.20±0.00 60.55±3.09
F1 50.34±0.01 41.29±0.00 40.09±5.01 48.17±0.38 45.17±2.27 53.21±0.53 40.21±0.00 51.08±2.06 56.05±0.99 56.02±0.00 57.18±2.09
Purity 69.86±0.00 63.87±0.00 64.63±2.20 71.88±0.20 71.17±1.99 76.37±0.47 68.25±0.00 69.80±1.66 77.06±0.62 71.32±0.00 77.19±2.79
BO ACC 60.58±0.13 58.54±1.01 47.23±4.27 58.68±0.01 58.48±1.12 45.95±3.74 60.53±0.00 72.70±3.18 62.18±3.45 66.10±0.00 77.66±1.28
$\mathcal{K}$ 57.40±0.15 54.93±1.12 42.68±4.63 55.06±0.01 55.09±1.22 41.46±4.06 57.19±0.00 70.40±3.48 59.14±3.73 63.17±0.00 75.79±1.39
NMI 69.19±0.01 74.23±0.10 64.14±1.88 67.12±0.18 68.47±1.34 62.40±2.86 68.01±0.00 76.02±2.15 73.48±1.10 71.57±0.00 82.90±1.30
ARI 50.08±0.03 52.91±0.63 38.51±4.18 48.08±0.39 44.85±1.26 35.46±3.93 43.76±0.00 58.70±3.66 54.06±2.87 48.15±0.00 69.03±2.38
Precision 58.51±0.07 53.43±1.93 43.22±7.31 60.02±3.43 57.02±1.38 41.68±4.58 67.13±0.00 72.81±4.06 61.06±4.54 67.66±0.00 73.90±2.42
Recall 62.54±0.13 55.47±1.56 46.23±3.84 55.30±0.03 60.88±1.24 47.71±3.40 60.50±0.00 70.82±3.34 61.44±3.81 66.81±0.00 77.65±2.39
F1 58.19±0.09 51.05±1.77 41.29±4.95 53.03±0.14 56.35±1.36 39.24±3.37 60.06±0.00 69.40±3.12 59.04±4.17 64.41±0.00 74.27±2.52
Purity 64.84±0.12 62.03±0.96 50.05±4.82 59.57±0.01 63.77±1.12 48.01±2.83 60.53±0.00 73.31±2.33 66.40±2.20 66.10±0.00 77.71±1.32
TR ACC 64.83±0.02 79.18±0.05 75.70±2.13 82.75±0.27 86.50±0.00 75.47±0.00 86.14±0.00 87.97±0.74 89.67±3.53 89.96±0.00 93.11±0.96
$\mathcal{K}$ 55.63±0.03 72.57±0.07 69.00±2.60 77.33±0.35 82.13±0.00 68.62±0.06 81.27±0.00 83.94±0.91 86.38±4.61 86.62±0.00 90.82±1.26
NMI 51.87±0.01 66.73±0.03 72.78±2.38 78.42±0.06 75.62±0.00 77.90±0.01 79.80±0.00 76.15±1.49 85.46±1.17 86.74±0.00 89.74±0.94
ARI 36.82±0.02 55.56±0.09 71.92±2.15 78.27±0.20 79.47±0.00 72.81±0.00 86.44±0.00 82.42±1.91 89.34±2.26 91.69±0.00 94.00±1.30
Precision 59.63±0.01 80.26±0.02 64.54±4.21 63.69±0.42 68.48±0.00 55.46±0.00 68.19±0.00 69.75±0.72 73.89±7.29 84.31±0.00 93.25±0.67
Recall 58.24±0.01 83.02±0.02 60.56±2.65 65.29±0.32 69.84±0.00 59.22±0.00 63.60±0.00 70.15±0.04 76.64±5.06 83.09±0.00 93.11±0.96
F1 55.25±0.02 78.78±0.03 61.57±3.22 60.65±0.49 68.75±0.00 55.81±0.00 62.83±0.00 69.73±0.44 74.31±6.71 82.88±0.00 92.61±0.95
Purity 65.61±0.01 79.18±0.05 87.27±1.68 82.94±0.27 87.09±0.00 86.87±0.00 88.08±0.00 88.66±0.81 91.75±2.81 89.96±0.00 93.66±0.74
@@
Figure 3: Clustering maps on Indian Pines: (a) ground truth, (b) K-means, (c) SSC, (d) NCSC, (e) SGLSC, (f) DGAE, (g) S3-ULDA, (h) S2DL, (i) SDST, (j) SPGCC, (k) SAPC, (l) Ours.
@@ IV-B Experiment Setup
2) Compared Methods: To thoroughly assess the clustering performance of the proposed method, we compare it against ten baseline methods, including K-means [12], SSC [14], NCSC [26], SGLSC [25], DGAE [33], S3-ULDA [42], S2DL [43], SDST [32], SPGCC [38] and SAPC [44]. For K-means, clustering is conducted at the pixel level. In contrast, SSC performs clustering at the superpixel level to mitigate the high memory cost. For other methods, we follow the settings in their papers and exhaustively search optimal values for hyperparameters. Here we briefly introduce each method:
@@
3) Evaluation Metrics: To quantitatively evaluate the clustering performance of different methods, we adopt eight mainstream clustering metrics: accuracy (ACC), kappa coefficient ($\mathcal{K}$), normalized mutual information (NMI), adjusted Rand index (ARI), Precision, Recall, F1-score (F1) and Purity. For Precision, Recall and F1, we use their macro-average form. Among these metrics, ACC, NMI, Precision, Recall, F1 and Purity range in [0, 1] while Kappa and ARI range in [-1, 1]. Each metric is reported as a percentage, with a higher value indicating the better clustering result. Following the common process, we adopt the Hungarian algorithm [45] to map the predicted labels based on the ground truth before calculating clustering metrics.
@@ IV-C Comparison with State-of-the-Art Methods
We report the clustering performance of our method (SSGCO-EGAEL) and other compared methods on Indian Pines, Pavia University, Botswana, and Trento in Table II. For each metric, the best value is marked in bold, while the second-best value is underlined.
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods with a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 90.9% (291/320) cases.
@@
Moreover, the clustering map visualization of different methods on Indian Pines is reported in Fig. 3. (Clustering map visualizations of other datasets are reported in the supplementary material.) According to the visualizations, it is evident that our method demonstrates superior clustering for large-scale and continuously distributed land covers, e.g., green and yellow areas in Indian Pines. This advantage primarily stems from our method performing superpixel-level clustering and adopting the edge learning to construct a more accurate adjacency matrix for superpixels.
@@
S3-ULDA     11.5 374.5 2792.4 88.9     816.8     9
S2DL     12.5 544.9 2960.6 65.7     895.9     10
SDST     296.9 229.7 169.0 129.2     206.2     6
SPGCC     85.7 619.9 1129.2 437.3     568.0     8

### v3

@@ I Introduction
The second limitation is the inaccurate superpixel topological graph. Typically, the superpixel topological graph is constructed based on either spatial adjacency or spectral similarity. The former connects spatially adjacent superpixels as neighbors [25, 38], while the latter determines neighbors based on spectral similarity [30, 32, 33]. EGRC-Net[31] adaptively refines the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings. However, it fails to take into account the reliability of existing edges. Since the superpixel topological graph is constructed without supervision, both spatial and spectral graphs inevitably contain incorrect edges that connect two superpixel nodes of different classes. Accordingly, these incorrect edges bring semantically different neighbors, leading to distorted node representations and confused class semantics during GNN aggregation. Fig. 1b illustrates the impact of deleting incorrect edges on clustering results. As the deletion ratio increases, clustering performance gradually improves, highlighting the importance of dealing with incorrect edges in the graph.
@@ III-D Training Procedure
We illustrate the training procedure of the proposed method in algorithm 1. In each epoch, we first compute the empirical edge weights and the predicted edge weights to update the adjacency matrix and the corresponding symmetrically normalized matrix. Second, we perform forward-propagation, computing node representations, prototypes and loss functions. Third, network parameters are optimized. Fourth, under evaluation mode, we apply spherical K-means on the output of $f^{\prime}(\cdot)$ to get the current superpixel clustering assignments and then update the prototypes. Finally, we transform the clustering result from the superpixel level to the pixel level by assigning each superpixel’s label to its internal pixels.
@@
5:  Update $\widehat{\mathbf{A}}$ based on the current $\mathbf{A}$ by Eq. (5).
6:  Randomly sample pixels to get $\{\bm{x}^{+}\}$.
7:  Forward $f(\bm{x}),f^{\prime}(\bm{x}^{+})$ by Eq. (12).
8:  Compute $\mathcal{L}_{NA}$ by Eq. (7).
9:  Compute $\{\bm{\mu}\},\{\bm{\mu}^{+}\}$ by Eq. (8).
10:  Compute $\mathcal{L}_{PC}$ by Eq. (9).
11:  Compute $\mathcal{L}$ by Eq. (22).
12:  Update $f(\cdot)$ with SGD optimizer.
13:  Update $f^{\prime}(\cdot)$ by Eq. (11).
14:  Evaluation mode, forward $f^{\prime}(\bm{x})$, perform K-means to get $\mathbf{R}$, and compute $\{\bm{\mu}\}$ by Eq. (8).
15:end for
16: Transform $\mathbf{R}$ from the superpixel level to the pixel level.
17:return $\mathbf{R}$
@@ IV-A Datasets
TABLE II: Clustering Performance on HSI Datasets, Indian Pines, Pavia University, Botswana and Trento are Abbreviated to IP, PU, BO and TR, $\bullet$ Indicates that our Method is Statistically Superior to the Compared Method According to the Pairwise $t$-Test at 0.05 Significance Level.
@@
Ours
IP     ACC     34.72±1.00 47.99±0.46 53.48±0.84 58.29±1.19 57.10±3.39 56.02±1.87 65.41±3.37 64.70±0.00 50.58±4.46 67.40±1.17 55.88±0.00 70.01±1.95
    $\mathcal{K}$     28.39±1.10 42.13±0.57 48.54±1.12 54.54±1.22 52.18±3.56 49.80±2.51 61.10±3.50 60.24±0.00 44.01±4.27 63.62±1.34 49.59±0.00 66.55±2.12
    NMI     43.24±0.57 57.81±0.44 59.47±1.40 63.65±0.35 63.68±1.03 56.91±1.69 67.65±1.65 71.97±0.00 50.21±1.80 67.54±1.26 54.03±0.00 71.17±1.31
    ARI     21.02±0.99 31.46±0.55 36.13±1.42 40.21±0.49 45.35±4.75 38.72±3.50 48.99±3.76 52.93±0.00 31.51±3.83 56.02±1.94 36.51±0.00 55.99±2.07
    Precision     37.74±1.15 43.93±1.72 42.74±2.88 51.88±1.62 50.59±1.81 42.50±4.87 51.34±1.87 59.63±0.00 45.34±2.71 52.66±2.19 49.61±0.00 58.72±2.23
    Recall     40.27±0.38 47.42±5.06 51.12±0.62 57.11±1.97 49.79±3.89 46.36±4.65 51.11±3.16 59.08±0.00 46.37±1.38 61.34±3.48 52.39±0.00 65.77±4.42
    F1     34.24±0.87 40.65±1.34 40.76±0.92 49.50±1.76 46.00±2.58 39.26±3.12 48.07±3.10 55.19±0.00 44.30±2.74 52.30±1.89 44.67±0.00 57.65±2.38
    Purity     52.26±0.87 57.78±0.51 62.87±0.65 69.58±1.25 68.94±1.51 59.78±2.92 72.52±2.06 75.32±0.00 56.51±1.36 73.94±1.78 59.33±0.00 77.18±1.03
PU     ACC     51.15±0.02 55.17±0.00 56.08±3.69 57.24±0.34 58.56±5.21 59.91±3.00 59.42±0.28 60.13±0.00 61.03±1.80 62.62±0.38 63.94±0.00 70.00±3.46
    $\mathcal{K}$     41.49±0.06 45.57±0.00 44.89±5.06 48.28±0.36 49.15±5.07 47.08±2.91 50.98±0.47 49.06±0.00 50.18±2.12 54.82±0.46 53.02±0.00 62.29±3.65
    NMI     53.97±0.01 51.48±0.00 47.34±4.27 53.69±0.39 53.72±2.58 54.59±3.71 62.93±0.08 48.15±0.00 56.41±1.44 57.84±1.20 61.89±0.00 61.90±3.18
    ARI     31.02±0.03 35.26±0.00 42.84±5.04 42.64±0.20 43.65±6.69 43.10±2.61 49.04±0.18 43.12±0.00 42.94±2.63 46.19±0.89 44.91±0.00 57.26±8.57
    Precision     51.83±0.01 43.00±0.00 40.91±4.93 50.29±0.39 48.16±4.44 46.73±5.54 56.77±0.70 48.68±0.00 53.97±3.11 56.32±1.16 56.21±0.00 57.82±1.22
    Recall     53.59±0.01 48.81±0.00 46.19±7.95 53.75±0.35 50.75±3.47 49.13±6.92 55.63±0.51 50.37±0.00 55.34±1.58 58.96±0.42 59.20±0.00 60.55±3.09
    F1     50.34±0.01 41.29±0.00 40.09±5.01 48.17±0.38 45.17±2.27 43.76±5.71 53.21±0.53 40.21±0.00 51.08±2.06 56.05±0.99 56.02±0.00 57.18±2.09
    Purity     69.86±0.00 63.87±0.00 64.63±2.20 71.88±0.20 71.17±1.99 68.39±2.92 76.37±0.47 68.25±0.00 69.80±1.66 77.06±0.62 71.32±0.00 77.19±2.79
BO     ACC     60.58±0.13 58.54±1.01 47.23±4.27 58.68±0.01 58.48±1.12 62.83±5.75 45.95±3.74 60.53±0.00 72.70±3.18 62.18±3.45 66.10±0.00 77.66±1.28
    $\mathcal{K}$     57.40±0.15 54.93±1.12 42.68±4.63 55.06±0.01 55.09±1.22 59.63±6.28 41.46±4.06 57.19±0.00 70.40±3.48 59.14±3.73 63.17±0.00 75.79±1.39
    NMI     69.19±0.01 74.23±0.10 64.14±1.88 67.12±0.18 68.47±1.34 73.76±3.93 62.40±2.86 68.01±0.00 76.02±2.15 73.48±1.10 71.57±0.00 82.90±1.30
    ARI     50.08±0.03 52.91±0.63 38.51±4.18 48.08±0.39 44.85±1.26 54.34±4.30 35.46±3.93 43.76±0.00 58.70±3.66 54.06±2.87 48.15±0.00 69.03±2.38
    Precision     58.51±0.07 53.43±1.93 43.22±7.31 60.02±3.43 57.02±1.38 53.14±9.53 41.68±4.58 67.13±0.00 72.81±4.06 61.06±4.54 67.66±0.00 73.90±2.42
    Recall     62.54±0.13 55.47±1.56 46.23±3.84 55.30±0.03 60.88±1.24 60.31±6.37 47.71±3.40 60.50±0.00 70.82±3.34 61.44±3.81 66.81±0.00 77.65±2.39
    F1     58.19±0.09 51.05±1.77 41.29±4.95 53.03±0.14 56.35±1.36 53.17±8.22 39.24±3.37 60.06±0.00 69.40±3.12 59.04±4.17 64.41±0.00 74.27±2.52
    Purity     64.84±0.12 62.03±0.96 50.05±4.82 59.57±0.01 63.77±1.12 64.05±6.06 48.01±2.83 60.53±0.00 73.31±2.33 66.40±2.20 66.10±0.00 77.71±1.32
TR     ACC     64.83±0.02 79.18±0.05 75.70±2.13 82.75±0.27 86.50±0.00 77.73±1.93 75.47±0.00 86.14±0.00 87.97±0.74 89.67±3.53 89.96±0.00 93.11±0.96
    $\mathcal{K}$     55.63±0.03 72.57±0.07 69.00±2.60 77.33±0.35 82.13±0.00 69.32±2.75 68.62±0.06 81.27±0.00 83.94±0.91 86.38±4.61 86.62±0.00 90.82±1.26
    NMI     51.87±0.01 66.73±0.03 72.78±2.38 78.42±0.06 75.62±0.00 75.54±0.81 77.90±0.01 79.80±0.00 76.15±1.49 85.46±1.17 86.74±0.00 89.74±0.94
    ARI     36.82±0.02 55.56±0.09 71.92±2.15 78.27±0.20 79.47±0.00 73.90±1.91 72.81±0.00 86.44±0.00 82.42±1.91 89.34±2.26 91.69±0.00 94.00±1.30
    Precision     59.63±0.01 80.26±0.02 64.54±4.21 63.69±0.42 68.48±0.00 51.10±2.14 55.46±0.00 68.19±0.00 69.75±0.72 73.89±7.29 84.31±0.00 93.25±0.67
    Recall     58.24±0.01 83.02±0.02 60.56±2.65 65.29±0.32 69.84±0.00 55.75±3.76 59.22±0.00 63.60±0.00 70.15±0.04 76.64±5.06 83.09±0.00 93.11±0.96
    F1     55.25±0.02 78.78±0.03 61.57±3.22 60.65±0.49 68.75±0.00 49.50±4.57 55.81±0.00 62.83±0.00 69.73±0.44 74.31±6.71 82.88±0.00 92.61±0.95
    Purity     65.61±0.01 79.18±0.05 87.27±1.68 82.94±0.27 87.09±0.00 77.86±2.00 86.87±0.00 88.08±0.00 88.66±0.81 91.75±2.81 89.96±0.00 93.66±0.74
@@
Figure 3: Clustering maps on Indian Pines: (a) ground truth, (b) K-means, (c) SSC, (d) NCSC, (e) SGLSC, (f) DGAE, (g) EGRC, (h) S3-ULDA, (i) S2DL, (j) SDST, (k) SPGCC, (l) SAPC, (m) Ours.
@@ IV-B Experiment Setup
2) Compared Methods: To thoroughly assess the clustering performance of the proposed method, we compare it against eleven baseline methods, including K-means [12], SSC [14], NCSC [26], SGLSC [25], DGAE [33], EGRC [31], S3-ULDA [42], S2DL [43], SDST [32], SPGCC [38] and SAPC [44]. For K-means, clustering is conducted at the pixel level. In contrast, SSC performs clustering at the superpixel level to mitigate the high memory cost. For other methods, we follow the settings in their papers and exhaustively search optimal values for hyperparameters. Here we briefly introduce each method:
@@
EGRC [31] refines the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings and is trained under the autoencoder framework.
@@
3) Evaluation Metrics: To quantitatively evaluate the clustering performance of different methods, we adopt eight mainstream clustering metrics: accuracy (ACC), kappa coefficient ($\mathcal{K}$), normalized mutual information (NMI), adjusted Rand index (ARI), Precision, Recall, F1-score (F1) and Purity. For Precision, Recall and F1, we use the macro-average form. Among these metrics, ACC, NMI, Precision, Recall, F1 and Purity range in [0, 1] while Kappa and ARI range in [-1, 1]. Each metric is reported as a percentage, with a higher value indicating the better clustering result. Following the common process, we adopt the Hungarian algorithm [45] to map the predicted labels based on the ground truth before calculating clustering metrics.
@@ IV-C Comparison with State-of-the-Art Methods
We report the clustering performance of our method and other compared methods on Indian Pines, Pavia University, Botswana, and Trento in Table II. For each metric, the best value is marked in bold, while the second-best value is underlined.
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods with a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 91.8% (323/352) cases.
@@
Moreover, the clustering map visualization of different methods on Indian Pines is reported in Fig. 3. (Visualizations of other datasets are reported in Supplement I.) According to the visualizations, it is evident that our method demonstrates superior clustering for large-scale and continuously distributed land covers, e.g., green and yellow areas in Indian Pines. This advantage primarily stems from our method performing superpixel-level clustering and adopting the edge learning to construct a more accurate adjacency matrix for superpixels.
@@
EGRC     18.5 40.7 324.9 42.1     106.5     5
S3-ULDA     11.5 374.5 2792.4 88.9     816.8     10
S2DL     12.5 544.9 2960.6 65.7     895.9     11
SDST     296.9 229.7 169.0 129.2     206.2     7
SPGCC     85.7 619.9 1129.2 437.3     568.0     9